# Hydra-NeXt: Robust Closed-Loop Driving with Open-Loop Training

### v1

@@ 2.3 Diffusion-based Driving Policies
Diffusion policies [8, 49, 6] have been used widely for the generation of robot behaviors. They prove effective in capturing multi-modal action distributions and generating smooth trajectories. In autonomous driving, recent works leverage diffusion models to predict trajectory waypoints given privileged input of surrounding traffic scenes [29, 17, 50, 47]. Our approach differs by employing the diffusion policy to generate multiple smooth and high-frequency control sequences, which serve as additional proposals for trajectory refinement.
@@ 3 Trajectory Decoder: Hydra-MDP
These metrics are used as rule-based teachers in Hydra-MDP. Fig. 2 shows the performance enhancement brought by rule-based heads, which appear to be limited compared with observations on NAVSIM [13]. This is likely caused by randomly disappearing agents in Bench2Drive  [^1] https://github.com/Thinklab-SJTU/Bench2Drive/issues/31, which confuse the rule-based heads. e.g. A trajectory that extends far into the future may be labeled as safe if its future waypoints do not intersect with a disappeared agent, even though the agent is currently visible and close to the ego vehicle. In the next section, we propose Hydra-Next, an extended version of Hydra-MDP addressing the main limitations in closed-loop driving.
@@ 4.1 Control Decoder $\pi_{ctrl}$
Our goal is to enable quick responses to reactive agents through direct control output, instead of relying on trajectory waypoints for rapid decisions. This is because waypoints are typically transformed into control signals through specific controllers such as a PID Controller or Model Predictive Control [5], which can lead to errors during path following. Further, TCP [48] points out that ensembling the trajectory and the control signal can boost driving performance, but it only considers a single modality for both the trajectory and the control prediction branches. This limitation makes it challenging to address uncertainties in environments [7] and makes it susceptible to interpolate between different action modes [24].
@@ 4.2 Trajectory Refinement $\pi_{dp}$
Another important aspect of closed-loop driving is to ensure the smoothness of the driving process. Benchmarks like nuPlan [4] and NAVSIM [13] post-process trajectories using an LQR tracker [43] and a kinematic bicycle model [41, 40] to adhere to kinematic constraints. Nevertheless, this idea has been neglected by existing open-loop E2E AD methods, which mainly focuses on cloning trajectory waypoints without considering kinematic constraints in closed-loop environments. This oversight becomes particularly critical when the vehicle experiences control loss in closed-loop environments [1, 27] or encounters large compounding errors [12] when quickly recovering to predicted waypoints is difficult.
@@ 5.2 Implementation Details
The implementation of Hydra-NeXt is largely consistent with open-loop E2E baselines [28, 23] on Bench2Drive. First, we train Hydra-NeXt on the training data for 20 epochs with a total batch size of 256, using 8 NVIDIA V100 GPUs. The AdamW [36] optimizer is used with the Cosine Annealing Scheduler [37] at a learning rate of $2\times 10^{-4}$ and a weight decay of 0.01. During training, data augmentations such as random cropping and photometric distortion are applied to the input images, which are first resized to a resolution of $800\times 450$.
@@ 5.3 Main Results
As shown in Tab. 1 and Tab. 2, Hydra-NeXt surpasses all E2E methods on Bench2Drive in key metrics such as the DS and SR distinctively. Specifically, under the CARLA v2 protocol, Hydra-NeXt outperforms the previous state-of-the-art, DriveAdapter [25], by 22.98 DS and 17.49% SR. Under the Bench2Drive protocol, it achieves improvements of 9.64 DS and 16.92 SR, despite DriveAdapter utilizing expert features from Think2Drive [31] for training. Additionally, Hydra-NeXt consistently demonstrates higher efficiency than all baselines, especially those using expert features. Nevertheless, Hydra-NeXt falls behind TCP [48] in terms of Comfort, which is likely due to frequent braking during interactive situations.
@@ 5.4 Ablation Study
Design of Trajectory Refinement. To illustrate the necessity of using the Diffusion Policy $\pi_{dp}$ as the proposal generator, we first replace the Diffusion Policy with a discrete control decoder $\pi_{ctrl}^{\prime}$ with a similar architecture to $\pi_{ctrl}$, which can also generate $N$ proposals based on top-$N$ confidence scores. This leads to a performance degradation of 5.78 DS as shown in Tab. 5, , which potentially stems from the poor ability of $\pi_{ctrl}^{\prime}$ to generate long and smooth control sequences. Further, we investigate the ability of $\pi_{dp}$ in modeling high-frequency control sequences. Tab. 6 indicates a performance degradation when using a low-frequency $\pi_{dp}$. On the contrary, a low-frequency $\pi_{ctrl}$ is better than a high-frequency one. This echos the previous finding and proves the effectiveness of employing $\pi_{dp}$ as the generator. Finally, Tab. 7 examines the impact of varying the number of proposals. Increasing the proposal number $N$ from 5 to 20 leads to saturating performance on DS, while SR mildly fluctuates with a larger $N$. The fluctuation may result from over-confident matching between $\pi_{dp}$ and the other decoders, while we can potentially benefit from small differences between their predictions for exploration.
@@ 7.1 $\pi_{traj}$: Trajectory Decoder
$\begin{array}[]{c}y_{i}=\frac{e^{-(\hat{T}-T_{i})^{2}}}{\sum_{j=1}^{k}e^{-(\hat{T}-T_{j})^{2}}}\\[1.42271pt] \mathcal{L}_{im}=-\sum_{i=1}^{k}y_{i}\log(\mathcal{S}^{im}_{i}).\end{array}\vspace{-0.15cm}$ (1)
@@
$\mathcal{L}_{traj}=\mathcal{L}_{im}+\mathcal{L}_{kd}.\vspace{-0.15cm}$ (3)
@@ 7.2 $\pi_{ctrl}$: Control Decoder
The loss of $\pi_{ctrl}$ computes a classification-based loss for each control signal (i.e. brake, throttle, and steer) across $t_{ctrl}$ timesteps. For simplicity, a control signal tuple $C=(brake,throttle,steer)$ is abbreviated as $(b,th,s)$, while the expert demonstration is denoted as $(\hat{b},\hat{th},\hat{s})$. Specifically, these loss functions are formulated as:
$\begin{array}[]{c}\mathcal{L}_{brake}=\sum_{t=1}^{t_{ctrl}}Focal(b^{t},\hat{b^{t}})\\[1.42271pt] \mathcal{L}_{throttle}=\sum_{t=1}^{t_{ctrl}}CE(th^{t},\hat{th^{t}})\\[1.42271pt] \mathcal{L}_{steer}=\sum_{t=1}^{t_{ctrl}}CE(s^{t},\hat{s^{t}})\\[1.42271pt] \mathcal{L}_{ctrl}=\mathcal{L}_{brake}+\mathcal{L}_{throttle}+\mathcal{L}_{steer},\end{array}\vspace{-0.15cm}$ (4)

### v2

@@ 2.3 Diffusion-based Driving Policies
Diffusion policies [8, 52, 6] have been used widely for generating robot behaviors. They prove effective in capturing multi-modal action distributions and generating smooth trajectories. In autonomous driving, recent works leverage diffusion models to predict trajectory waypoints given privileged input of surrounding traffic scenes [30, 17, 53, 50]. Our approach differs by employing the diffusion policy to generate multiple smooth and high-frequency control sequences, which serve as additional proposals for trajectory refinement.
@@ 3 Trajectory Decoder: Hydra-MDP
These metrics are used as rule-based teachers in Hydra-MDP. Fig. 2 shows the performance enhancement brought by rule-based heads, which appear to be limited compared with observations on NAVSIM [13]. This is likely caused by randomly disappearing agents in Bench2Drive  [^1] https://github.com/Thinklab-SJTU/Bench2Drive/issues/31, which confuse the rule-based heads. e.g. A trajectory that extends far into the future may be labeled as safe if its future waypoints do not intersect with a disappeared agent, even though the agent is currently visible and close to the ego vehicle. In Sec. 4, we propose Hydra-NeXt, an extended version of Hydra-MDP addressing the limitations in closed-loop driving.
@@ 4.1 Control Decoder $\pi_{ctrl}$
Our goal is to enable quick responses to reactive agents through direct control output, instead of relying on trajectory waypoints for rapid decisions. This is because waypoints are typically transformed into control signals through specific controllers such as a PID Controller or Model Predictive Control [5], which can lead to errors during path following. Further, TCP [51] points out that ensembling the trajectory and the control signal can boost driving performance, but it only considers a single modality for both the trajectory and the control prediction branches. This limitation makes it challenging to address uncertainties in environments [7] and susceptible to interpolate between different action modes [25].
@@ 4.2 Trajectory Refinement $\pi_{dp}$
Another important aspect of closed-loop driving is to ensure the smoothness of driving. Benchmarks like nuPlan [4] and NAVSIM [13] post-process trajectories using an LQR tracker [46] and a kinematic bicycle model [42, 41] to adhere to kinematic constraints. Nevertheless, this idea has been neglected by existing open-loop E2E AD methods, which mainly focuses on cloning trajectory waypoints without considering kinematic constraints in closed-loop environments. This oversight becomes particularly critical when the vehicle experiences control loss in closed-loop environments [1, 28] or encounters large compounding errors [12] when quickly recovering to predicted waypoints is difficult.
@@ 5.2 Implementation Details
The implementation of Hydra-NeXt is largely consistent with open-loop E2E baselines [29, 23] on Bench2Drive. First, we train Hydra-NeXt on the training data for 20 epochs with a total batch size of 256, using 8 NVIDIA V100 GPUs. The AdamW [37] optimizer is used with the Cosine Annealing Scheduler [38] at a learning rate of $2\times 10^{-4}$ and a weight decay of 0.01. During training, data augmentations such as random cropping and photometric distortion are applied to the input images, which are first resized to 800×450800$\times$ 450800 × 450.
@@ 5.3 Main Results
As shown in Tab. 1 and Tab. 2, Hydra-NeXt surpasses all E2E methods on Bench2Drive in key metrics such as the DS and SR distinctively. Specifically, under the CARLA v2 protocol, Hydra-NeXt outperforms the previous state-of-the-art, DriveAdapter [26], by 22.98 DS and 17.49% SR. Under the Bench2Drive protocol, it achieves improvements of 9.64 DS and 16.92 SR, despite DriveAdapter utilizing expert features from Think2Drive [32]. Additionally, Hydra-NeXt demonstrates higher efficiency than the baselines, especially those using expert features. Nevertheless, Hydra-NeXt falls behind TCP [51] in terms of Comfort, which is likely due to frequent braking during interactive situations.
@@ 5.4 Ablation Study
[…]
@@ 6 Conclusion
Acknowledgement This work was supported in part by the National Natural Science Foundation of China (Grant 62472098). This work was supported by the Science and Technology Commission of Shanghai Municipality (No. 24511103100).
@@ 7.1 $\pi_{traj}$: Trajectory Decoder
$\begin{array}[]{c}y_{i}=\frac{e^{-(\hat{T}-T_{i})^{2}}}{\sum_{j=1}^{k}e^{-(\hat{T}-T_{j})^{2}}}\\[1.42271pt] \mathcal{L}_{im}=-\sum_{i=1}^{k}y_{i}\log(\mathcal{S}^{im}_{i}).\end{array}\vskip-4.26773pt$ (1)
@@
$\mathcal{L}_{traj}=\mathcal{L}_{im}+\mathcal{L}_{kd}.\vskip-4.26773pt$ (3)
@@ 7.2 $\pi_{ctrl}$: Control Decoder
[…]
$\begin{array}[]{c}\mathcal{L}_{brake}=\sum_{t=1}^{t_{ctrl}}Focal(b^{t},\hat{b^{t}})\\[1.42271pt] \mathcal{L}_{throttle}=\sum_{t=1}^{t_{ctrl}}CE(th^{t},\hat{th^{t}})\\[1.42271pt] \mathcal{L}_{steer}=\sum_{t=1}^{t_{ctrl}}CE(s^{t},\hat{s^{t}})\\[1.42271pt] \mathcal{L}_{ctrl}=\mathcal{L}_{brake}+\mathcal{L}_{throttle}+\mathcal{L}_{steer},\end{array}\vskip-4.26773pt$ (4)
@@ 8 Performance of Individual Policies
9 Efficient Diffusion Policy $\pi_{dp}$
Method Latency (ms) $\downarrow$ D.S. $\uparrow$ S.R. $\uparrow$
VAD 224.3 39.42 10.00
Hydra-NeXt (w/o $\pi_{dp}$) 250.6 60.40 48.10
Hydra-NeXt* (DDPM) 528.3 65.89 48.20
Hydra-NeXt* (DDIM) 243.9 64.87 46.63
Table 11: Efficiency of Different Diffusion Schedulers. * denotes Flash-attention [11].
We found that the efficiency of $\pi_{dp}$ can be greatly enhanced without sacrificing too much performance. By replacing the DDPM with a DDIM scheduler [45], we can reduce the latency by approximately 53%, decreasing the number of denoising steps from 100 to 20. Additionally, by combining this modification with flash attention, we achieve latencies comparable to VAD while maintaining strong closed-loop performance, with only a marginal 1% drop in the Driving Score. Therefore, we conclude that $\pi_{dp}$ can be both effective and efficient with the right design choices.
@@ 10 Implementation on NAVSIM
11 Implementation on CARLA-Garage
Method Expert D.S. $\uparrow$ S.R. $\uparrow$
Hydra-NeXt Think2Drive 73.86 50.00
TF++† [24] PDM-Lite 84.21 67.27
SimLingo [44] PDM-Lite 85.94 66.82
Hydra-NeXt*† PDM-Lite 86.00 68.18
Table 12: Performance of Models Trained with Different Experts. †Ensemble of three models trained with different seeds.
We provide additional results on the CARLA-Garage dataset [24] collected by the PDM-Lite expert [2]. Compared to the Bench2Drive dataset [28], this dataset features smooth driving with no jittery behavior. We experiment with $\pi_{traj}$ and $\pi_{dp}$ within the TF++ framework [24]. Specifically, we replace the longitudinal control head of TF++ with $\pi_{traj}$ (imitation + collision head) and fuse the outputs of $\pi_{dp}$ with longitudinal and lateral controls. We then ensemble three models trained with different seeds. This variant, Hydra-NeXt*, achieves a Driving Score of 86%, surpassing existing methods.